I do not know where to start with this problem:
A compass and a ruler together cost $4. The compass costs .90 more than the ruler. How much does the compass cost?
I am 53 and have never understood how to do word problems. I can solve the equations but cannot come up with the equation itself.

Found 2 solutions by scott8148, dragonwalker:
Answer by scott8148(6628)

"A compass" ___ c

"and a ruler" ___ r

"together cost $4" ___ c + r = $4

"The compass costs .90 more than the ruler" ___ c = r + $.90

Answer by dragonwalker(73)

The key to coming up with a formula is to look at the first unknown and then the relationship between the unknowns.
We don't know the cost of either but we DO know that one costs more than the other and we know the difference.
We want to know how much the compass costs so let us give that value the letter 'a'
So if the compass costs 'a' then the ruler costs 0.90 less which can be simply written as a - 0.90
We know that the cost of both is $4.00 so:
we add 'a' (the cost of the compass) to a - 0.90 (the cost of the ruler).
So:
a + a - 0.90 = 4.00
rearrange:
a + a = 4.00 + 0.90
2a = 4.90
a = 2.45
So the cost of the compass is $2.45
If you want to check you can find from this that the ruler costs 90c less so costs $1.55
Add them together:
1.55 + 2.45 = 4.00 !!!!
Hope this helps...